mark is three times as old as jane is now. in three years, mark will be twice as old as jane is then,. how old are they now
Answer by stanbon(75887)

mark is three times as old as jane is now.
m = 3j
-----
in three years, mark will be twice as old as jane is then.
(m+3) = 2(j+3)
----
how old are they now
---
Substitute for "m" and solve for "j":
3j+3 = 2j+6
j = 3 (Jane's age now)
---
Since m=3j, m = 9 (Mark's age now)
====================================
Cheers,
Stan H.
============